OpenStudy (anonymous):

Help!!!! Determine whether the following integrals converge, and if they do, to what. Integral from 4 to infinity (1/x^3) dx

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \int\limits_{4}^{n} \frac{ 1 }{ x ^{3} }\]

OpenStudy (anonymous):

does that help or do you need more...?

OpenStudy (anonymous):



OpenStudy (anonymous):

i don't know how to do it

OpenStudy (anonymous):

evaluate the integral =-1/(2x^2) plug in n and 4 -1/(2n^2) - -1/(2(4)^2) find the limit as n-> infinity (hint: this one is super easy)

OpenStudy (anonymous):

clear on that?

OpenStudy (anonymous):

questions? comments? concerns?

OpenStudy (anonymous):

I need to see you do it out please?

OpenStudy (anonymous):

i got -1/32 ...but i think its suppose to be positive?

OpenStudy (anonymous):

- - 1/(2(4)^2)

OpenStudy (anonymous):

why is there 2 minuses?

OpenStudy (anonymous):

n -1/(2x^2) | = -1/(2n^2) - -1/(2*4^2) 4

OpenStudy (anonymous):

4 is the lower limit of integration, n is the upper limit of integration